Problem 6 Calculate the test statistic. (A z-score and a t-score are examples of test statistics.) In order to present your final results with a numerical accuracy of 0.01, it will be necessary to perform your intermediate calculations with an accuracy of 0.001 Part 1 (6%) The cost of a daily newspaper varies from city to city. A study was done to test the claim that the average cost of a daily newspaper is $0.50 Ten cities were surveyed; the average cost was $0.50 with a standard deviation of $0.06 Part 2 (6%) You statistics instructor claims that 60 percent of the students who take an Elementary Statistics class go through ife feeling more enriched. some mysterious reason, most people dont believe this. You decide to check this out on your own. You randomly survey 50 former Elementary Statistics students and find that 35 feel more enriched as a result of the cless.
Answer #1

Answer

Part 1

we have μ 0.50. 0.50. s = 0.06, n = 10

mu is population mean, ar{x} is sample mean, n is sample size and s is standard deviation.

Using the formula t = (ar{x}-mu)/(s/sqrt{n})

setting the given values, we get

t=(0.5 - 0.5)/(0.06/V10) 0/(0.06/V10) 0 ..............................(0 divided by anything gives us 0)

So, calculated t value is 0.00

Part 2

we have population proportion p = 60% = 60/10 = 0.60

sample proportion hat{p} = 35/50 = 0.70

sample size is n =50

Formula for z statistic is given as

z = (hat{p}-p_o)/sqrt{(p_o*(1-p_o))/n}

setting the given values, we get

z = (0.70-0.6)/V/(0.6 * (1-0.60)/50 0.10/0.069 1.44

So, test statistic is z = 1.44